OpenStudy (anonymous):

I have to integrate cos^2(x). Wolfram tells me that I will have to write cos^2(x) as (cos2(x)+1)/2 to proceed. Why is that? Isn't there a direct method of integrating trigonometric functions that have powers? If yes, why? :)

OpenStudy (amistre64):

potatoe potatoe

OpenStudy (anonymous):

That didn't help.

OpenStudy (amistre64):

its more effective when read in context :) its 2 different ways of getting to the same results

OpenStudy (amistre64):

i say potatoe, you say potatoe ....

OpenStudy (amistre64):

the reduction formula is the "direct" method

OpenStudy (amistre64):

lets see if i can recall it correctly:\[\int cos^n=\frac{1}{n}cos^{n-1}sin+\frac{n-1}{n}\int cos^{n-2}\]

OpenStudy (anonymous):

Which method is more intuitive? I have opened an MIT lecture about power reduction on another tab, by the way.

OpenStudy (amistre64):

cos^2 is small enough that rewriting it in its identity tends to be simpler than trying to recall the formula

OpenStudy (amistre64):

the intergral of cos^n is by parts; whereas cos^2 is just an identity